OpenStudy (anonymous):

Can someone help me please? For the given quadratic equation convert into vertex form, find the vertex, and find the value for x = 6. y = -2x2 + 2x +2

OpenStudy (anonymous):

Do you know how to complete the square?

OpenStudy (anonymous):

no I do not

OpenStudy (anonymous):

You should look it up. Give an effort to learn it. I'll help if you try and it doesn't work.

OpenStudy (anonymous):

btw that is -2x^2

OpenStudy (anonymous):

I know

OpenStudy (anonymous):

every example so far they haven't had to substitute 6 for x. Not only that but they don't have a negative x^2 and they = 0 not y.

OpenStudy (anonymous):

I really like the effort that you have put forth.

OpenStudy (anonymous):

\[ y = -2x^2 + 2x +2 \]If you divide by \(-2\), then: \[ -\frac{y}{2} = x^2-x-1 \]

OpenStudy (anonymous):

Do you think you can complete the square on the right side now?

OpenStudy (anonymous):

Don't worry about substitute for \(6\) just yet.

OpenStudy (anonymous):

I am really lost on what the next step would be. From what I see you would try to remove the square but I don't know how to do that.

OpenStudy (anonymous):

Now, you create the square.

OpenStudy (anonymous):

A square would be the following:\[ (x+a)^2 \]

OpenStudy (anonymous):

Where \(a\) is some number.

OpenStudy (anonymous):

When you foil, you get: \[ (x+a)^2=x^2+2ax+a^2 \]

OpenStudy (anonymous):

We have to match it up with what we have.

OpenStudy (anonymous):

Since we have \(x^2 + (-1)x+(-1)\)

OpenStudy (anonymous):

The \(2a =-1\) when we match it up.

OpenStudy (anonymous):

This means \(a=-1/2\)

OpenStudy (anonymous):

And \(a^2=1/4\).

OpenStudy (anonymous):

So \[ (x-1/2)^2 = x^2-x+1/4 \]

OpenStudy (anonymous):

This would mean that: \[ x^2-x -1 + 1/4 - 1/4 = (x-1/2)^2-1-1/4 = (x-1/2)^2-5/4 \]

OpenStudy (anonymous):

\[ -y/2= (x-1/2)^2-5/4 \]We can multiply by negative \(-2\) again and: \[ y=-2(x-1/2)^2+5/2 \]

OpenStudy (anonymous):

This is the vertex form.

OpenStudy (anonymous):

I am so confused

OpenStudy (anonymous):

Okay, I will do it one more time.

OpenStudy (anonymous):

I really don't understand how you went from (x+a)^2 to all of that.

OpenStudy (anonymous):

We have \(x^2-x\), we only want to focus on this part.

OpenStudy (anonymous):

We take the coefficient of \(x\), in this case it is \(-1\), because \(x^2-x = x^2+(-1)x\)

OpenStudy (anonymous):

Do you follow?

OpenStudy (anonymous):

No